OpenStudy (anonymous):

FACTORING BY GROUP: 1. y^2 - 8y - y + 8 2. 25 - (x^2-6x+9) 3. x^2 + 4xy + 4y^2 - 64

Parth (parthkohli):

1) Factorize the first two terms and last two terms, and try to factorize even further. 2) First, distribute the negative sign. Then, find such an arrangement that you are able to factorize the first and last two terms with the same in-bracket thing. Now see 1. 3) First two and last two :D

OpenStudy (anonymous):

1)y^2 - 8y - y + 8=y^2-9y+8=(y-1)(y-8)

OpenStudy (anonymous):

\[ \frac{ y ^{2} }{ y ^{} }-\frac{ 8y }{ y }-(y-8)\]

OpenStudy (anonymous):

OK I got it

OpenStudy (anonymous):

Let me try the others

OpenStudy (anonymous):

2. 25 - (x^2-6x+9)=25-x^2+6x-9=-x^2+6x+16=-(x^2-6x-16)=-(x-8)(x+2)=(8-x)(x+2)

OpenStudy (anonymous):

3. x^2 + 4xy + 4y^2 - 64=(x+2y)^2-64=(x+2y)^2-8^2=(x+2y+8)(x+2y-8) hope it helps!

OpenStudy (anonymous):

I'm confused about problem 2 and 3

OpenStudy (anonymous):

which step?

OpenStudy (anonymous):

Lets begin with problem 2. I did as Looking for trinomials. SO I got (5-x+3)(5+x-3)

OpenStudy (anonymous):

I'm still confused. Looking at examples and trying to figure out. Also #3. I don't get it

OpenStudy (ajprincess):

(5-x+3)(5+x-3) 5-x+8-x 5+x-3=2+x (5-x+3)(5+x-3)=(8-x)(2-x) You get the same answer as @mymathcourses even if u do in a different way bcause that is the simplified answer.:) Getting it @rosimary?

OpenStudy (anonymous):

Yes. I'm understanding

OpenStudy (anonymous):

I was confused because the sample the teacher gave for that particular problem she stopped at (4-x+y)(4+x-y). I guess she is teaching us by step? I'm not sure but I'm understanding.

OpenStudy (anonymous):

I am still confused as to what to do about problem 3. I got a weird answer that looks like (x+1)(x+4y)(2y-8)(2y+8)

OpenStudy (anonymous):

yes , it should use a^2- b^2=(a+b)(a-b)

OpenStudy (anonymous):

a^2 year is (x+Y)^2 and b^2 here is 8^2

OpenStudy (anonymous):

which problem are you ref too?